the altitude of an equilateral triangle is 4.2 meters long find the perimeter of the triangle
Answer by TimothyLamb(4379)

s = sqrt( 4.2*4.2 + (s/2)(s/2) )
s = sqrt( 4.2*4.2 + (ss/4) )
ss = 4.2*4.2 + (ss/4)
ss - (ss/4) = 4.2*4.2
(4ss/4) - (ss/4) = 4.2*4.2
(3ss/4) = 4.2*4.2
(3/4)ss = 4.2*4.2
ss = 4.2*4.2*(4/3)
s = sqrt( 4.2*4.2*(4/3) )
s = 2*4.2 * sqrt( 1/3 )
s = 4.84974226119
---
check:
sqrt( (4.84974226119/2)*(4.84974226119/2) + 4.2*4.2 ) = 4.84974226119
---
answer:
perimeter = 3 * 4.84974226119
perimeter = 14.5492267836 meters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php